calculate the horizontal component of the initial velocity of the shot when the shot is a metal ball of mass 4.0 kg and the angle of projection for a particular throw is 40 degree
1
Expert's answer
2020-03-18T10:17:30-0400

The horizontal component of the initial velocity is


"v_x=v_0\cdot \cos 40°" .


Assuming that "v_0=5" "m\/s"


We have


"v_x=v_0\cdot \cos 40°=5\cdot \cos 40°=3.83" "m\/s"